OpenStudy (anonymous):

In 2005, about 5 billion gal renewable fuel, such as ethanol and other biofuels, were in the united states. The energy law passed in 2007 requires that 40 billion gal such fuels be used by 2022. What is the percent of increase?

OpenStudy (anonymous):

5 billion in 2005 number of years between 2007 and 2022 = 2022 - 2007 = 15 40/15 = 2.666 which is a decrease of 53% i think the badly worded question wants to know the increase from 5 to 40 think of it like this 5 x percentage = 40 so percentage = 40/5 = 8 so 800%

OpenStudy (anonymous):

thanks